Two trains are running in opposite directions with same speed. If the length of each train is 135 m and they cross each other in 18 sec. Find the speed of each train.

Grade:9

1 Answers

Arun
5 years ago
Distance covered = 135 + 135 = 270 m

Time = 18 s

Let the speed of each train = v. Then relative speed = v + v = 2v

2v = distance/time = 270/18 = 15 m/s

Speed of each train = v = 15/2 = 7.5 m/s 

= 7.5 × 36/10 km/hr = 27 km/hr
 
 
Regards
Arun (askIITians forum expert)